Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, boldrini, mantovani, ruggi - 17:40 Thursday 23 September 2021 (53247) Print this report
ISC Shift: lock acquisition adaptation after RH turn on

Today we worked on the recovery and adaptation of the lock acquisition after the turning on of the Ring Heaters on the ETMs:

  • the first thing we noticed is an increased value of B1p_DC when the DRMI is locked (CARM offset = 3 kHz), which is now ~ 0.024 mW (previously it was ~ 0.015 mW);
  • the DRMI seems quite easier to lock, even keeping the SRM aligned; more statistics is needed to confirm this;
  • we had a lot of unlocks of the West ALS, either when engaging the boost filters on the beating signals (which causes an unexpected dip in the green transmitted power, which then slowly recovers), or in random places of the lock acquisition, in this cases with a faster but deeper glitch;
  • throughout the lock acquisition, we had an overall reduction of the MICH gain (and, to a lesser extent, PRCL gain);
  • we had a couple of unlocks due to DARM_CORR being cut to zero; we debugged the logic and we caught the two places where the effect originated; we added a new manual switch, LSC_GREEN_OFF, which is set to 1 once DARM is moved back to the IR beam and the automatic unlock safety is active; such switch prevents the unlock of the ALS from unlocking everything; this was tested by manually unlocking the West ALS at such point of the lock acquisition, with no effect on the ITF;
  • we went a few times up to 480 mW of arm powers (which now correspond to a bigger CARM offset), re-checking gains and demodulation phases; unfortunately we could not stay for a long time in such place, as usually an oscillation at 55 Hz well visible on DARM would kill the lock (e.g. 13:00:09 UTC); we reduced DARM gain but the investigation on this oscillation will continue in the next shifts; also, a retuning of the B1p demodulation phases (both PD and QPD at 56 MHz) could be needed;
  • we went once to CARM Null, where we could stay locked for a bit more than 4 minutes (from around 12:32:20 UTC); for this first lock we decided not to add the following steps (SSFS, additional ASC loops, 1f handoffs) and to study the behaviour of the ITF:
    • the power transmitted from the arm cavities has increased to ~ 850 mW as maximum peak, then it decreased to around ~ 750 mW;
    • the value of the dark fringe power is much lower and much more stable than before (Figure 1);
    • the sidebands power are decreasing more quickly over time.

The retuning and optimization of the ITF will continue in the afternoon shift.

Images attached to this report
Search Help